OpenStudy (silverlark2):

The gradient of the line joining the points (9,a) and (2a,1) is 2/a where a is not equal to 0. Calculate the possible values of a. Ans: 3 or -6 but I can't get the answer.

OpenStudy (anonymous):

\[\frac{a-1}{9-2a}=\frac{2}{a}\] is what you need to solve

OpenStudy (anonymous):

\[a(a-1)=2(9-2a)\] \[a^2-a=18-4a\] \[a^2+3a-18=0\] etc

OpenStudy (anonymous):

\[3^2+18=0\]

OpenStudy (anonymous):

18^2 has a A

OpenStudy (anonymous):

so solve for a